1. Find the domain: square root of 10-2x
2. Simplify without radicals square root of (25x^2 + 16)^2
3. Write as a radical expression A^2/5
4. Rewrite under one radical square root of x times square root of (x^4)^5

Thank You Very Much!!!

Answer by stanbon(57967)

1. Find the domain: square root of 10-2x
10-2x >= 0
2x<=10
x<=5
Domain: all Real numbers less than or equal to five
--------------------------
2. Simplify without radicals square root of (25x^2 + 16)^2
25x^2+16
--------------------
3. Write as a radical expression A^2/5
sqrt(A^4/25)
----------------------
4. Rewrite under one radical square root of x times square root of (x^4)^5
= sqrt(x*x^20)
= sqrt(x^21)
========================
Cheers,
Stan H.